OpenStudy (princesssleelee):

A ball is thrown from an initial height of 4 feet with an initial upward velocity of 21/fts. The ball's height h (in feet) after t seconds is given by the following. h=4+21t-16t^2 Find all values of t for which the ball's height is 10 feet.Round your answer(s) to the nearest hundredth. (If there is more than one answer, use the "or" button.)

OpenStudy (princesssleelee):



OpenStudy (kinged):

Equate the equation to 10, and solve for t.

OpenStudy (princesssleelee):

what do you mean 'equate'? @Kinged

OpenStudy (oops):

he means set it equal to

OpenStudy (kinged):

Make it equal to 10.

OpenStudy (princesssleelee):

10=4+21t-16t^2

OpenStudy (kinged):

Yep.

OpenStudy (princesssleelee):

t= 0.89, 0.42

OpenStudy (kinged):

Nope. Use quadratic formula and take care of the signs.

OpenStudy (kinged):

You'll find two answers: one positive and one negative. Dismiss the negative and take the positive to court.

OpenStudy (princesssleelee):

is it going to be a fraction with a square root in it? @Kinged

OpenStudy (kinged):

Just show your answer.

OpenStudy (princesssleelee):

\[t=\frac{ 21\pm \sqrt{57} }{32 }\]

OpenStudy (kinged):

Your answer should be in decimal.

OpenStudy (kinged):

Again, you'll find two answers: one positive and one negative. Dismiss the negative and take the positive to court.

OpenStudy (princesssleelee):

I have no idea....

OpenStudy (kinged):

Evaluate the expression with a plus sign, then evaluate it again with a minus sign.

OpenStudy (kinged):

That's what the plus/minus sign means.

OpenStudy (princesssleelee):

10=4+21t-16t^2 -10=""

OpenStudy (princesssleelee):

0.89, -0.48

OpenStudy (kinged):

Use this: http://www.math.com/students/calculators/source/quadratic.htm

OpenStudy (kinged):

Substitute the values accordingly and that's about it.

OpenStudy (princesssleelee):

i have no idea what that grid was asking for me to fill in , im equaling to 10 as they are 0

OpenStudy (kinged):

What's the value before t^2? That's A. What's before t? That's B. What's before C? That's the number.

OpenStudy (princesssleelee):

-0.23, -1.08

OpenStudy (kinged):

Nope. Did you substitute the right values in A, B, and C?

OpenStudy (princesssleelee):

well, i thought i did...

OpenStudy (kinged):

-16 goes in A, 21 goes in B, and 4 goes in C.

OpenStudy (princesssleelee):

i forgot the negative sign for 16

OpenStudy (princesssleelee):

-0.16, 1.48

OpenStudy (kinged):

Hey! That's it. Dismiss the negative value, now. That is it.

OpenStudy (princesssleelee):

0.16, 1.48

OpenStudy (kinged):

I clearly typed "dismiss the negative value" not "dismiss the negative sign".

OpenStudy (princesssleelee):

so only 1.48

OpenStudy (kinged):

Yeah.